The forecasted cash flows from a stock investment are $6 each year for the next 70 years. The first dividend is one year from today. After that, the company will cease to exist and have no value. What is the per share value of this stock today if the required rate of return is 10%?

Please show formula and process used
Answer #1

Per share value of stock today is $ 59.92

As per dividend discount model, current share price is the present value of future dividends.
Present value of dividends = Annual dividend * Present value of annuity of 1
= $       6.00 * 9.987338
= $    59.92
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.10)^-70)/0.10 i = 10%
= 9.987338 n = 70